OpenStudy (anonymous):

MEDAL AND FAN Sylvia has a total of 16 trading cards: 8 superhero cards, 4 princess cards, 2 villain cards, and 2 wildlife cards. What is the ratio of wildlife cards to the total number of trading cards in simplest form? ? to ?

OpenStudy (danjs):

What is the total of all the cards?

OpenStudy (danjs):

and What is the number of wildlife cards?

OpenStudy (alexandervonhumboldt2):

2:16=2/16=1/8=1:8

OpenStudy (science_ally):

2:16 in simpliest form is 1:8

OpenStudy (anonymous):

so 1 to 8?

OpenStudy (anonymous):

yes

OpenStudy (science_ally):

yep

OpenStudy (science_ally):

so 9:36 in simpliest form is 1:4